### v3

# Scientific Literature Hierarchy Construction and Organization

Abstract
Scientific knowledge is growing rapidly, making it difficult to track progress and high-level conceptual links across broad disciplines. While tools like citation networks and search engines help retrieve related papers, they lack the abstraction needed to capture the density and structure of activity across subfields.
We motivate Science Hierarchography, the goal of organizing scientific literature into a high-quality hierarchical structure that spans multiple levels of abstraction—from broad domains to specific studies. Such a representation can provide insights into which fields are well-explored and which are under-explored. To achieve this goal, we develop a hybrid approach that combines efficient embedding-based clustering with LLM-based prompting, striking a balance between scalability and semantic precision. Compared to LLM-heavy methods like iterative tree construction, our approach achieves superior quality-speed trade-offs. Our hierarchies capture different dimensions of research contributions, reflecting the interdisciplinary and multifaceted nature of modern science. We evaluate its utility by measuring how effectively an LLM-based agent can navigate the hierarchy to locate target papers. Results show that our method improves interpretability and offers an alternative pathway for exploring scientific literature beyond traditional search methods. 111 Code, data and demo are available: https://github.com/JHU-CLSP/science-hierarchography
Science Hierarchography:
Hierarchical Organization of Science Literature
Muhan Gao, Jash Shah, Weiqi Wang, Daniel Khashabi Department of Computer Science Johns Hopkins University Baltimore, MD 21218, USA {mgao38,jshah48,wwang194,danielk}@jhu.edu
1 Introduction
The pace of scientific publishing is accelerating (Ware and Mabe, 2015), but this growth is uneven across fields (Hope et al., 2023). Some areas attract dense research activity, while others remain underexplored. This raises a natural question:
How do we understand the distribution of scientific efforts across different sub-areas?
Answering this question is essential for both academic and policy stakeholders. A clearer view of how research efforts are distributed enables institutions to spot emerging or neglected areas, prioritize strategic hiring and future agendas. For policymakers, it supports more informed funding decisions, ensuring that critical but underexplored domains receive the attention and resources they deserve.
Figure 1: An example of  Science Hierarchography illustrates how scholarly work can be organized hierarchically—from broad research domains at the top, through increasingly specific sub-clusters, down to individual papers at the lowest level. Critically, this structure must be inferred automatically and at scale.
Conventional tools like Google Scholar are designed as retrieval engines, optimized to return a handful of papers that match a specific query. They offer little in the way of a comprehensive or structured view of the broader scientific landscape. Similarly, while modern LLM-based assistants can surface related works (seen during pretraining or via their retrieval tools), they fall short in offering a broad, bird’s-eye perspective on scientific progress.
Addressing this challenge requires abstraction: a way to generalize over research problems and techniques and to connect broad scientific areas to specific papers via intermediate categories. At one end, we have high-level domains (e.g., physics, AI); at the other, individual papers. Between them lie a latent spectrum of subfields and methodological clusters. What’s missing is a data structure that captures all these abstraction levels.
We propose building large-scale hierarchical representations of scientific literature, which we call Science Hierarchography. A well-designed hierarchy provides a macro-level view of scientific progress, revealing how research is distributed across methods and application areas. This helps researchers spot emerging trends and gaps, and supports policymakers and institutions in making more strategic resource decisions. It also offers a new way to explore the literature—complementing traditional search by allowing users to navigate science through conceptual hierarchies.
How should scholarly work be represented? A central challenge in building a scientific hierarchy is defining what each node represents. Research papers often span multiple topics (e.g., reinforcement learning for medical imaging or deep learning for oceanography). To capture this complexity, we develop a prompting strategy that decomposes papers into key contribution types—such as the problems addressed and techniques used3.2). For each fixed contribution type, we construct a corresponding hierarchical structure, ensuring that papers are organized into meaningful, coherent categories.
What construction strategies balance scalability and quality? To address this, we introduce Scychic (pronounced “psychic”), a new method for building high-quality hierarchical structures of scientific literature. Scychic integrates fast embedding-based clustering with LLM prompting, combining the efficiency of embeddings with the semantic precision of language models (§4.1).
How can we evaluate the quality of a scientific hierarchy? Scientific hierarchies lack a fixed ground truth—they evolve over time as research landscapes shift. We therefore adopt an evaluation-through-utilization approach, measuring whether an information seeker (human or AI) can efficiently locate specific content (e.g., child nodes) by navigating the hierarchy from the root. This evaluation hinges on the idea that a good hierarchy enables rapid information discovery, even though its utility extends well beyond search alone (§5.2).
What did our empirical results show? Our approach achieves the best trade-off between quality and speed when compared to LLM-heavy methods like iterative tree construction or pruning. Extensive experiments show that Scychic consistently produces higher-quality hierarchies than a broad set of baselines (§5.4). Validation on a 10K-paper dataset further confirms its strong accuracy and scalability for large-scale use.
Contributions: (1) We introduce the goal of constructing large-scale, abstract hierarchies of scientific literature to reveal how scholarly efforts are distributed across research areas. (2) We propose a utilization-based evaluation framework that measures how effectively users can discover information by traversing the hierarchy. (3) We present Scychic, a new method that combines fast embedding-based clustering with LLM prompting to build high-quality, multidimensional hierarchies. Extensive experiments show that Scychic outperforms baseline approaches, offering a more structured and bird’s-eye view of scientific progress.
System # of Levels Node content Node granularity Assigned by Purpose Public
Web of Science One Research areas One keyword Publisher Indexing No
Scopus Two Research areas One keyword Editor Indexing Yes
arXiv Taxonomy Two Research areas One keyword Authors Indexing Yes
PubMed MeSH Multiple Medical headings One keyword Authors Indexing Yes
Microsoft Academic Graph Multiple Research areas Multiple keywords Algorithms Indexing Discontinued
Science Hierarchography (Ours) Multiple (by designer) Rich contribution descriptions Science contribution summary (many tokens) Algorithms Exploratory Analysis Yes
Table 1: Comparison of hierarchical resources for organizing scientific literature, ordered by hierarchy depth. Conventional systems are built for indexing, relying on fixed, shallow taxonomies with keyword-based nodes and human-assigned labels. In contrast, Science Hierarchography supports deeper, designer-controlled hierarchies with rich natural-language summaries, enabling more flexible and exploratory analysis of scientific work.
2 Related Work
Gierarchy induction: The field of taxonomy induction has progressed from early pattern-based techniques to modern LLM-augmented methods. Seminal work by Hearst (1992) introduced the use of hand-crafted hyponym patterns for extracting is-a relationships. Subsequent research expanded on this using statistical methods and large-scale information extraction to identify hypernym-hyponym structures (Pantel and Pennacchiotti, 2006; Yang and Callan, 2009; Girju et al., 2006).
Recent advances incorporate LLMs prompting to enhance taxonomy construction. For example, Wan et al. (2024); Zeng et al. (2024a); Chen et al. (2023); Zeng et al. (2024b) apply zero-/few-shot reasoning and ensemble ranking, while others explore open-ended, vocabulary-free taxonomy creation (Gunn et al., 2024), self-supervised expansion in low-resource domains (Mishra et al., 2024), and graph-based methods leveraging metadata and citations (Cong et al., 2024; Sas and Capiluppi, 2024; Shen et al., 2024). Optimization and in-context learning have also shown promise (Hu et al., 2024b; Shi et al., 2024; Xu et al., 2025; Jain and Espinosa Anke, 2022; Chen et al., 2021).
Our work differs in scope, scale, and methodological design. We focus on scaling taxonomy induction for the domain of scholarly literature—a setting that presents greater challenges than typical setups (e.g., entity hierarchy) due to the complexity, size, and evolving nature of scientific content. The most comparable effort is by Oarga et al. (2024), though our broader objectives require fundamentally different algorithmic strategies and operate without access to ground truth labels.
Structured representation of science: As science grows at an unprecedented rate (Teufel et al., 1999; Pertsas and Constantopoulos, 2017; Constantin et al., 2016; Fisas et al., 2016; Liakata et al., 2010), numerous frameworks have emerged to structure this information through knowledge graphs and taxonomies (Fathalla et al., 2017; Jaradeh et al., 2019; Oelen et al., 2020; Vogt et al., 2020; Soldatova and King, 2006). Recent work includes prompt-based topic modeling (Pham et al., 2024), iterative taxonomy construction that incorporates object properties and graph mining (Cui et al., 2024; Marchenko and Dvoichenkov, 2024), and hybrid approaches that combine curated ontologies with data-driven maps (Zimmermann et al., 2024). Our work builds on these efforts by constructing a high-quality hierarchical structure tailored to scientific literature, in three key ways. The prior work: (1) Produces shallow hierarchies, typically only one or two levels deep; (2) Uses cluster labels based on keywords, whereas ours are derived from natural language summaries of papers; (3) Depends heavily on manual effort, while our pipeline is fully automated.
In Table 1 we summarize the differences with existing hierarchical resources. While most prior systems are limited to one/two level(s) of depth and rely on manually assigned labels for indexing—a process often prone to bias (Hadfield, 2020)—our approach supports deeper, algorithmically generated hierarchies with semantically rich node descriptions. This enables a more flexible and interpretable representation of scientific knowledge.
3 Science Hierarchography: Toward Hierarchy of Scholarly Work
We begin with a formal problem definition (§3.1), followed by content representation (§3.2) and depth considerations (§3.3).
3.1 Formal Problem Statement
We define the task of Science Hierarchography as an inference problem where the input is a large set of scientific papers: $P=\{p_{1},p_{2},\ldots,p_{n}\}$. The goal is to infer a hierarchical structure for these papers (e.g., their problems and techniques). This structure represents levels of specificity and abstraction, with nodes closer to the root representing broader topics. Broader topics are at the upper levels, while more specific subtopics and individual papers are at the lower levels.
Formally, this goal is defined as inferring a graph that encodes relationships between nodes (atomic concepts representing scholarly ideas or goals). This is formally defined as $T\vcentcolon=(N,E)$ where:
  • NN is the set of all concepts (nodes) present in the hierarchy: $N=\{n_{1},n_{2},\ldots,n_{N}\}$.
  • $E$ is the set of all edges connecting node pairs: $E=\{r_{1},r_{2},\ldots,e_{M}\}$ where each edge is a node pair $r_{i}=(n_{p},n_{c})$ connects a parent node $n_{p}$ to its child node $n_{c}$.
  • For each edge $e\in E$, there is a type attribute: $\text{type}(e)\in\{\text{``isA''},\text{``instanceOf''}\}$ associated with the edge.
  • Each node $n\in N$ has a property $n.prop$ which defines the content of the node (e.g., a sentence description or a few keywords).
The edge (relations connecting two nodes) have types that is either “isA” relationships or “instanceOf.” The "isA" relationship defines a hierarchical link between node pairs, indicating a child node is a subclass of its more abstract parent node (e.g., "RLHF isA RL" means "RLHF" is a type of "RL"). The "instanceOf" relationship indicates that a specific instance belongs to a concept category (e.g., "Ouyang et al. (2022) instanceOf RLHF" means "Ouyang et al. (2022)" is a specific instance of the broader “RLHF” cluster).
3.2 Decomposing Papers to Contributions
A central challenge is how to represent the content of scholarly work within hierarchy nodes. Scientific papers are idea-dense, often combining broad goals, specific problems, and technical methods. To capture this complexity, we extract structured representations that disentangle these distinct aspects (D’Souza and Auer, 2020). This also mitigates the issue of input length: papers typically range from 4 to 10 pages (5K to 10K tokens), making full-document processing across large corpora infeasible and costly for LLMs.
We use an LLM to preprocess each paper (title and abstract) and break them down into a pre-defined set of contributions, akin to prior work Hope et al. (2017); Chan et al. (2018) that mine “problem schema” from existing documents. We consider the following contribution types: (1) problem statement (the problem addressed), (2) solution (the technical approach used), (3) result (the key finding), and (4) topic (the overarching themes). (See §C for prompts and examples). We note that each contribution may include additional dimensions (sub-contributions). For instance, a “result” encompasses both the “outcome” and its “potential impact.”
3.3 Choosing Hierarchy Depth
While the ideal number of hierarchy layers is ultimately empirical, we can build useful intuition from the structure of a near-balanced tree. For a tree with branching factor $b$ and depth $L$, the total number of nodes is roughly $O(b^{L})$. To organize $C$ contributions, the number of nodes should scale with $C$, implying a depth of $L=O(\log_{b}C)$. In practice, we use $L=3$ for a 2K-paper corpus and $L=4$ for 10K papers, consistent with this logarithmic scaling. Extrapolating further, corpora of $10^{7}$ papers would likely require depths of $L=6$ or $7$.
4 Tackling Science Hierarchography
We present algorithms to address our proposed goal. We start with our main method, Scychic4.1), explore its special cases (§4.2), and then describe alternative baselines that rely more heavily on LLMs (fLMSci; §4.3). While all approaches leverage LLMs to some extent, they differ significantly in their reliance on them: some require many calls (linear or quadratic in the number of papers), while others are more efficient (e.g., logarithmic). Since our goal is to scale to millions of papers, minimizing LLM usage is critical. Our objective is to identify the method that yields the highest-quality hierarchy with the lowest LLM overhead, balancing quality, latency, and cost.
4.1 Scychic: Alternating Between Clustering and Summarization
This approach is based on the following design choices: (1) access to embedder, a neural model that converts a description into a $d$-dimensional vector, (ideally) capturing its semantic meaning; (2) a clustering algorithm clusterer that, given the hyperparameter $k$, generates $k$ clusters; (3) a contribution type (e.g., problem definition) and its dimensions $C$ extracted per paper as detailed in §3.2 which determines the focus of the node descriptions; (4) summarizer, an LLM that generates a summary description which (ideally) provides a more abstract description of a collection of node descriptions; and (5) the total number of hierarchy layers $L$ and target number of clusters in each layer $(k_{1},k_{2},\ldots,k_{L})$.
Initialization: The approach begins by embedding each papers. For each paper $p_{i}$, we embed each component in $C$: $\texttt{{\color[rgb]{0.09765625,0.0078125,0.6796875}embedder}}(c_{j}^{i})\in\mathbb{R}^{d}$, where $j\in C$. This process results in $|C|$ embeddings per paper. We concatenate these embeddings, yielding $\mathbb{R}^{d.|C|}$ embeddings per paper. With this, we present the main algorithm which involves two phases:
Top-down phase: First, we use a top-down strategy, iterating through the top half of the layers. For each layer $l\in(1\ldots\left\lfloor L/2\right\rfloor)$ (starting from the root and moving downward), we cluster the documents using clusterer.
For the first layer ($L=1$), we partition the entire paper collection into $k_{1}$ top-level clusters using their embeddings. For the second level, we process each top-level cluster independently, i.e., within each top-level cluster, we take only the papers belonging to that cluster and apply clusterer to their embeddings to divide them into subgroups.
This process continues recursively: for subsequent layers, clustering operates on the data points within each cluster. We run as many clustering algorithms as there are clusters in the previous layer. The allocation of the total cluster count across different top-level clusters is determined proportionally based on the number of papers in each top-level cluster. If a top-level cluster contains a larger proportion of papers, it will be allocated more subclusters from the total cluster count. This proportional allocation ensures that areas with higher paper density receive finer-grained clustering.
This recursively subdivides each cluster at level $l$ into multiple clusters at level $l+1$. This top-down phase continues until we reach level $\left\lfloor L/2\right\rfloor$. Alongside clustering, we also use the summarizer to generate abstracted summaries for each of the clusters. The generated cluster description follows the same structure or style as the input descriptions. For example, if the inputs are statements about problem categories, the output from summarizer is also in the same style, but more abstract.
Bottom-up phase: In the second phase, we switch to a bottom-up strategy to construct the remaining levels ($\left\lfloor L/2\right\rfloor+1$ through $L$). To form clusters for bottom-level (layer $L$), we apply clusterer to the paper embeddings within each sub-cluster within level-$\left\lfloor L/2\right\rfloor$ (the lowest level clustering obtained from top-down approach). We then use the summarizer to create an abstracted description for each cluster.
We repeat this process for all layers from $L$ to $\left\lfloor L/2\right\rfloor+1$. To build layer $l$, we start by embedding the generated cluster summaries from the level below l1l-1 using embedder, similar to how we embedded the papers. We then run the clustering clusterer on these new embeddings and generate abstracted summaries for the clusters to group these summaries into higher-level clusters. This bottom-up aggregation continues until we connect with the previously constructed level $\left\lfloor L/2\right\rfloor$ clusters.
Algorithm 1 Scychic algorithm
1:Set of papers $P=\{p_{1},p_{2},\ldots,p_{n}\}$, embedder, clusterer, summarizer, num of layers $L$, target cluster sizes $(k_{1},k_{2},\ldots,k_{L})$
2:Initialization: For each paper $p_{i}\in P$, embed their selected components to form $\mathbb{R}^{d\times|C|}$.
3:for layer $l=1$ to $\lfloor L/2\rfloor$ do $\triangleright$ Top-down phase
4:  if $l=1$ then
5:   Apply clusterer to divide papers into $k_{1}$ clusters
6:  else
7:   for each cluster from layer l1l-1 do
8:     Apply clusterer to divide into subclusters
9:  Use summarizer to generate summaries for each cluster
10:for each cluster $\kappa$ at level $\lfloor L/2\rfloor$ do $\triangleright$ Bottom-up phase
11:  for layer $l=L$ to $\lfloor L/2\rfloor+1$ do
12:   if $l=L$ then
13:     Collect the embeddings of papers within $\kappa$.
14:   else
15:     Apply embedder on summaries of cluster $l+1$
16:   Apply clusterer to form higher-level clusters
17:   Use summarizer to generate abstracted summaries
18:return Hierarchical structure
Rationale behind the hybrid design: The hybrid approach merges the strengths of top-down and bottom-up strategies. A bottom-up method may create less coherent top-level clusters. The top-down approach ensures high-quality top-level clusters but doesn’t utilize the abstracted summaries from summarizer used by bottom-up clustering. By combining both methods, the hybrid design achieves robust and effective clustering. Our empirical results in §5.4 demonstrate this approach’s strength by balancing quality and scalability.
4.2 Top-down and Bottom-up Baselines
We examine two special cases of Scychic: one using only a top-down strategy and the other solely with a bottom-up approach. These variants help isolate and evaluate the strengths and limitations of each method. Results are discussed in §5.4.
4.3 Pure LLM-based Baselines
We introduce baselines that heavily utilize LLM calls, based on the hypothesis that LLMs can make high-quality local decisions, collectively forming a robust global structure. The potential cost here is the need to make many LLM calls. We refer to these baselines as fLMSci (pronounced “flimsy”) and present two variants below.
Initializing a Seed Hierarchy: The first step involves creating a seed hierarchy, starting with the hierarchy of sciences from the Wikipedia page on branches of science.222en.wikipedia.org/wiki/Branches_of_science We made several adjustments to this hierarchy, detailed in §D, where the resulting seed hierarchy is also included.
fLMSci (parallel): Adding Many Items in Parallel with Few Prompts: This approach involves adding items to the seed hierarchy in parallel. First, gather all unique items extracted from different papers and create batches of 100 items each (batching because all of these items would not fit within the context window of LLM). Then, implement a multi-threaded program where each thread adds a batch of keywords to a clone of the seed hierarchy. Finally, merge (with a Python program, not an LLM call) all these cloned trees into a single comprehensive structure.
fLMSci (incremental): Expand the tree by incrementally adding items one at a time: This approach involves an iterative, layer-by-layer prompt that navigates the tree starting from the root node and performs specific actions. By default, the supported actions include: (a) Go down: select one of the nodes to traverse down to a lower layer in the tree; (b) Add sibling: Insert a new node at the same level; (c) Make parent: Create a new parent node; (d) Discard: Ignore the item, if no suitable location is found. This prompt are shown in Fig.11.
The available actions vary depending on the tree’s current position. To prevent placing detailed items too high in the tree (above layer 3), we hide the actions that involve nodes (b) and (c). This restriction provides more control and reduces errors, as observed in pilot experiments. If the traversal reaches a child node, we hide the option to move further down (a). Another issue identified was frequent mistakes in earlier layers, likely due to broad labels where multiple categories could fit. To address this, we provide more context for the first layer by replacing label strings with descriptive definitions, as shown in Fig.10.
4.4 Computational Complexity of Approaches
A major scalability bottleneck in hierarchy construction is the number of LLM calls. Let $C$ be the number of contributions (§3.2), $b$ the branching factor, and $L=O(\log_{b}C)$ the maximum depth for a near-balanced tree (§3.3). Our proposed algorithm, Scychic, requires $O\left(C/b\right)$ LLM calls for both its top-down and bottom-up variants. Among the LLM-based baselines discussed in §4.3, fLMSci (parallel) makes $O(C/l)$ calls (with $l$ as batch size), offering lower complexity but at the cost of reduced quality. In contrast, fLMSci (incremental) achieves higher accuracy but requires $O(C\log_{b}C)$ LLM calls due to root-to-leaf traversals during insertion. Empirically, the difference in LLM usage is significant: in our 2K-paper setup, fLMSci (incremental) makes 61K calls compared to just 322 for Scychic (subsection 5.3).
Approach # of LLM calls
Scychic $O\left(C/b\right)$
fLMSci (parallel) $O(C/l)$
fLMSci (incremental) $O(C\log_{b}C)$
Table 2: Computational complexity of hierarchy construction methods measured by LLM calls, with $C$ = contributions, $b$ = branching factor, and $l$ = batch size.
5 Experimental Setup and Results
We describe our experimental setup, including the diverse paper collection used for our experiments (§5.1) and the evaluation framework (§5.2).
5.1 Collection of Science Papers
We compile a collection of scientific papers spanning domains such as computer science, neuroscience, biology, oceanography, and their interdisciplinary intersections. Our initial analysis focuses on a smaller set of approximately $2K$ papers (referred to as SciPile), allowing for rapid iteration over design choices and assessment of scalability. We then extend our analysis to a larger collection of $10K$ papers, referred to as SciPileLarge. Details on data collection and filtering are provided in §F.
5.2 Evaluation as Utilization
Ideally, hierarchy quality would be evaluated against a gold standard—but no such reference exists, and scientific literature continually evolves. As a result, we adopt an evaluation framework based on utilization, independent of fixed ground truth.
We assess hierarchy quality by measuring how well it supports navigation and content discovery. Specifically, we use an LLM-based agent to locate target papers via tree traversal, tracking accuracy at each level and across the full hierarchy. A stronger hierarchy should better capture conceptual relationships and improve information-seeking efficiency. While our evaluation focuses on retrieval, the hierarchy’s utility extends beyond that.
Two key design choices guide our evaluation: (a) selecting appropriate queries, and (b) choosing a reliable LLM. For (a), we sample paper titles and abstracts as queries. Although we considered generating language questions from papers, pilot studies showed both approaches yield similar results, so we use the simpler method. For (b), we use Qwen2.5-32b-instruct, which demonstrated performance closest to GPT-4 among open models in our pilot evaluations (§B).
Starting at the top of the hierarchy, we provide the evaluator (LLM) with the query and the list of cluster descriptions (prompt in Fig. 2). The model must select the cluster it deems most relevant. We move to the next level only if the chosen cluster contains the target paper. At each subsequent level, the same query is used, but only subclusters of the previously selected (and correct) cluster are shown. This continues until the model reaches the level containing the target paper. We report two metrics: Strict Accuracy (Strict-Acc), which captures how often the model selects the exact correct node, and Layer-1 Accuracy (L1-Acc), which measures how often it selects the correct top-level subtree.
5.3 Experiment Design
We conduct few sets of experiments to evaluate our method (Scychic, §4.1) against the baselines (§4.2, 4.3) following the proposed evaluation protocol. Hyperparameter details for Scychic are provided in §G. The experiments are structured as follows: (1) We begin by comparing all methods on “topic” contributions, the simplest contribution type (Table 5.3). Due to the high computational cost, LLM-based baselines are evaluated only in this setting. (2) We then evaluate performance on more complex contributions (problem, solution, and results) using both SciPile and SciPileLarge to test scalability (Table 4). The results tables also report LLM Cost (average number of input tokens and LLM calls) and Hierarchy Structure (depth and branching factor per node).
Method Strict-Acc (%) $\uparrow$ L1-Acc (%) $\uparrow$ # of Calls $\downarrow$
Topic contributions
Scychic 14.9 $\pm$ 2.7 65.7 $\pm$ 4.4 322
[…]
Top-down
14.5 $\pm$ 4.7 62.5 $\pm$ 7.4 322
[…]
Bottom-up
13.9 $\pm$ 5.3 54.4 $\pm$ 12.7 322
lightgrayblack
[…]
fLMSci (par)
4.0 $\pm$ 2.8 32.0 $\pm$ 6.3 226
[…]
fLMSci (inc)
18.0 $\pm$ 5.3 91.0 $\pm$ 4.0 61K
[…]
Method Accuracy (%) LLM Cost Hierarchy Structure
Strict-Acc $\uparrow$ L1-Acc $\uparrow$
Avg. # of
Input Tokens $\downarrow$
# of
Calls $\downarrow$
Depth
Avg. Branching
Factor
Max. Branching
Factor
Dataset: SciPile (2K papers)
Contributions type: Problem Statement
Scychic 51.1 $\pm$ 3.8 81.7 $\pm$ 2.6 2624 20
[…]
Top-down
49.0 $\pm$ 3.7 80.3 $\pm$ 2.7 2953 322 3 7.1 18
[…]
Bottom-up
45.9 $\pm$ 5.0 69.3 $\pm$ 8.1 2177 16
Contributions type: Solution Statement
Scychic 48.8 $\pm$ 6.1 82.3 $\pm$ 1.1 2343 16
[…]
Top-down
45.9 $\pm$ 5.5 79.2 $\pm$ 3.4 2521 322 3 7.1 19
[…]
Bottom-up
36.7 $\pm$ 2.6 67.0 $\pm$ 4.3 1990 14
Contributions type: Results Statement
Scychic 46.4 $\pm$ 5.2 76.4 $\pm$ 6.9 2654 16
[…]
Top-down
47.3 $\pm$ 3.1 80.5 $\pm$ 4.4 2718 322 3 7.1 16
[…]
Bottom-up
40.0 $\pm$ 10.7 64.0 $\pm$ 8.9 2210 13
Dataset: SciPileLarge (10K papers)
Contributions type: Problem Statement
Scychic 43.7 $\pm$ 6.5 85.8 $\pm$ 4.2 7451 26
[…]
Top-down
41.5 $\pm$ 8.2 86.5 $\pm$ 5.6 8990 1572 4 8 30
[…]
Bottom-up
26.2 $\pm$ 5.4 41.9 $\pm$ 4.0 5924 26
Contributions type: Solution Statement
Scychic 24.7 $\pm$ 4.8 65.8 $\pm$ 2.5 7653 28
[…]
Top-down
22.4 $\pm$ 3.5 52.3 $\pm$ 3.0 4032 1572 4 8 26
[…]
Bottom-up
23.9 $\pm$ 3.3 51.3 $\pm$ 3.1 6150 28
Contributions type: Results Statement
Scychic 27.6 $\pm$ 4.6 69.8 $\pm$ 2.1 6457 30
[…]
Top-down
19.7 $\pm$ 4.0 54.0 $\pm$ 3.3 5380 1572 4 8 30
[…]
Bottom-up
23.6 $\pm$ 2.7 55.2 $\pm$ 2.9 4731 28
Table 4: Evaluation results of Scychic and the corresponding baselines on both the 2K (SciPile) and 10K (SciPileLarge) datasets. Scychic maintains high accuracy and a relatively small variance, proving the rationale behind our hybrid design. When scaling from 2K to 10K papers, our method shows a slight decrease in Strict-Acc but maintains strong L1-Acc, demonstrating its feasibility on larger datasets. Across both scales, the problem statement contribution type consistently yields the most accurate hierarchies, indicating this contribution type contributes most for hierarchy construction.
5.4 Empirical Findings
Scychic outperforms its special-case baselines. As shown in Table 5.3, Scychic achieves higher Level-1 accuracy than the top-down and bottom-up baselines, while maintaining comparable Strict-Acc. Similar trends hold across other contribution types in Table 4. For example, on “solution” contributions, Scychic exceeds the top-down baseline by 2.9% in Strict-Acc and 3.1% in L1-Acc, highlighting its effectiveness. Notably, these gains are achieved with a similar number of tokens and LLM calls, underscoring Scychic’s compute efficiency.
LLM-based baselines can be far more expensive than Scychic. While fLMSci slightly outperforms Scychic in accuracy, it does so at the cost of a massive increase in LLM calls—making it impractical for large-scale use. As a result, despite its strong performance, fLMSci (incremental) simply doesn’t scale.
Scychic scales to larger paper corpus. For our scalability experiments, we evaluated Scychic on our larger $10K$ paper dataset SciPileLarge, using the problem statement contribution type. Due to the significant increase ($\times 5$) in corpus size, we implemented a 4-layer hierarchy instead of the 3-layer structure used previously. Notably, Scychic achieved even higher L1-Acc (86.5%) on SciPileLarge compared to our smaller dataset SciPile. This improvement likely stems from the enhanced quality of our expanded dataset, which has more strict filtering mechanisms. While the Strict-Acc showed a minor decrease compared to results on SciPile, it remained at a satisfactory level. Collectively, these results provide compelling evidence that our method scales successfully to substantially larger paper corpora.
5.5 Additional Analyses
We briefly cover additional analyses that were omitted from the main text due to space constraints.
Detailed prompts significantly improve hierarchy quality. To demonstrate this, we compare two prompt types. The first is a "detailed" prompt—carefully curated with comprehensive instructions and reminders—which we use for all main experiments in this paper. The second is a "simplified" prompt containing only the core task description. The results confirm that the detailed prompt consistently and substantially outperforms the simplified version across all scenarios. More detailed results are in §H.3.
Embedding quality varies significantly across models. For the embedder mentioned in §4.1. We evaluated three models—Qwen’s gte-Qwen2-7B-instruct (Li et al., 2023), OpenAI’s text-embedding-3-large, and text-embedding-ada-002. The first two models perform similarly, whereas text-embedding-ada-002 produces markedly weaker results. We select gte-Qwen2-7B-instruct for its strong balance of performance and its practical value as an open-weight model for reproducible research. The experimental results are in §H.2.
5.6 Sample Visualization of the Hierarchy
The reader might be curious to see the resulting hierarchies. In §I we show a slice of the final hierarchy generated by Scychic on the SciPileLarge  dataset. The original hierarchy has 4 levels, use papers’ problem contribution. Due to space constraints, this slice shows only two levels of clusters above the individual papers.
6 Discussion and Conclusion
Future applications:
Our work opens several promising directions for future research. One key opportunity is to use the constructed hierarchies as tools for exploratory analysis across scientific domains. They can aid academic institutions and funding bodies in identifying emerging trends and underexplored areas, and can be adapted for domain-specific analyses that capture the unique structure of individual fields. This approach not only deepens our understanding of scientific progress but also provides a new lens for organizing the vast and growing body of scholarly work.
Conclusions:
We introduced Science Hierarchography, a framework for large-scale hierarchical summarization of scientific literature, offering a new lens on how research efforts are distributed. Our method, Scychic, combines LLMs with efficient algorithms to strike a balance between quality and scalability. Looking forward, we aim for this work to help researchers navigate the scientific landscape more intuitively and support more informed resource allocation in academia.
Limitations
Although we evaluated our pipeline on $10K$ papers, this is still far from the true scale of scientific literature. We hope future work will enhance our approach to handle more realistic scales. Additionally, while our evaluation framework shows potential for efficient information discovery, it may have its own weaknesses and biases. Integrating human verification into the assessment process could help ensure the quality and reliability of the inferred hierarchies.
Ethics Statement
In our work, all data and models are accessed via licenses that grant us free and open access for research purposes. Expert annotations are provided by the paper’s authors, who have contributed their efforts without compensation. We have not observed any harmful content in either the scholarly papers or the content generated by LLMs. On the other hand, since our resulting hierarchy reflects the distribution of scientific efforts across various fields, it offers a detailed map of where research activity is concentrated and where it is lacking. This nuanced view can guide decision-makers—such as government agencies and academic institutions—in making more informed choices about resource allocation. By highlighting underexplored yet promising areas alongside well-established fields, the hierarchy helps ensure that funding, support, and strategic initiatives are distributed more equitably. Ultimately, this balanced approach can foster innovation and drive progress in areas that might otherwise be overlooked, leading to a more inclusive and socially beneficial advancement of science.
Acknowledgments
This work is supported by ONR grant (N0001424-1-2089) and Defense Advance Research Projects Agency (DARPA) under Contract No. HR001125C0304. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of DARPA. We sincerely thank Jiefu Ou and the broader JHU CLSP community for discussions and inspiration.
Appendix A Additional Related Work
We include additional related work here because of the space limitation in the main text.
Clustering with LLMs: Recent advances in clustering methodologies augmented by LLMs have demonstrated effective ways to generate interpretable groupings of text. For example, Viswanathan et al. (2024); Katz et al. (2024) apply few-shot clustering and thematic grouping to partition scientific literature into meaningful subtopics, while Zhang et al. (2023); Wang et al. (2023) further refine these techniques by aligning clustering outcomes with natural language explanations and user intent. Other recent work iteratively refines cluster representations by replacing cluster centroids or summary points with LLM-generated natural language descriptions and inclusion criteria, thereby inducing more abstract, interpretable concepts over multiple clustering rounds (Lam et al., 2024; Diaz-Rodriguez, 2025). While these approaches improve clustering quality by using LLMs at various stages, they mostly result in flat groupings rather than hierarchical structures. Our approach builds on this by using LLMs to cluster documents and organizing these clusters into a structured hierarchy.
Structured knowledge in LLMs:
Prior work has explored how LLMs internalize hierarchical knowledge. For example, He et al. (2024); Lovón-Melgarejo et al. (2023); Park et al. (2025) extend the linear representation hypothesis to reveal that LLMs encode categorical concepts as polytopes, with hierarchical relationships reflected as orthogonal directions. Other works such as Wolfman et al. (2024) and Budagam et al. (2024) examine the benefits of explicit hierarchical syntactic structures and prompting frameworks for guiding LLM performance, while Moskvoretskii et al. (2024) and Hu et al. (2024a) focus on constructing and materializing large-scale structured knowledge bases about entities and events. In line with the same aspirations, our work explores the use of hierarchical structures to organize scientific literature.
Structured knowledge representation:
Understanding and organizing knowledge is a fundamental pursuit in both artificial and human intelligence (Dahlberg, 1993). Abstraction hierarchies, such as WordNet for lexical semantics (Miller, 1995), ConceptNet for commonsense reasoning (Speer et al., 2017), and Probase for large-scale concept representation (Wu et al., 2012), have proven to be powerful tools for structuring information. Similarly, modern tabular reasoning leverages structured representations to facilitate systematic inference and knowledge retrieval, demonstrating that such structure remains crucial (Wang et al., 2024).
Appendix B Evaluation Framework
We provide more context on our evaluation. As discussed in §5.2, we use randomly-sampled papers (title/abstract) as a query. The evaluator LLM goes through the hierarchy, starting from the root node and iteratively selects the relevant nodes to traverse. The prompt for each decision is shown in Fig.2.
Figure 2: Prompt used for Evaluation
One question is, which LLM should we use for evaluation? We chose Qwen2.5-32b-instruct for its strong instruction-following capabilities. In pilot experiments, Qwen showed a high consistency against GPT4 score, compared to other open-weight models. Here’s a summary of that experiment: We evaluated one of the hierarchies produced by Scychic using different models, including GPT-4. Assuming GPT-4 has the highest accuracy, we sought alternative models with the greatest consistency against it, as frequent evaluations with GPT4 are costly. Fig.5 presents the results. As it can be observed, Llama has the highest agreement, but we suspect bias since the hierarchy was also constructed with Llama. To avoid this, we selected the next best model, Qwen2.5-32b-instruct, for evaluation.
Evaluator LLLM Agreement with GPT4
GPT-3.5 39.6
GPT4-mini 59.2
Gemma3-24b-it 62.1
Qwen2.5-32b-instruct 66.5
Llama 3.3 70B 72.4
Table 5: Agreement of different evaluator LLMs against GPT4.
Appendix C Extracting Paper Contributions
Below are prompts and examples for extracting different contributions (problem, solution, result and topics) from papers’ titles and abstracts. we utilize the GPT-4o model (gpt-4o-2024-08-06) to generate all contribution extractions along with detailed rationales explaining the extraction decisions.
C.1 Prompt for Extracting Problem/Solution/Result Contributions
We use the prompt below to extract contributions from the paper’s title and abstract. After finishing the extraction, the three contributions will be saved into the original json file. Please see §3.2 for more information.
Figure 3: Prompt for extracting Problem/Solution/Result contributions
C.2 Prompt for Extracting Topic Contributions and Rationales
This section has the prompt of generating topics and rationales from papers given their titles and abstracts. The prompt provides the model with a system role instruction that describes the task, title, and abstract, and also an example to get the specified output format.
Figure 4: Prompt of Topic and Rationale Generation
C.3 Examples for Problem/Solution/Result/Topic contributions extracted from papers
Below we show examples of paper titles and abstracts, and different contributions (Problem/Solution/Result/Topic) we extract by language model.
Problem/Solution/Result/Topic contributions from scientific papers
Title: Sixfold excitations in electrides
Abstract: Due to the lack of full rotational symmetry in condensed matter physics, solids exhibit new excitations beyond Dirac and Weyl fermions, of which the sixfold excitations have attracted considerable interest owing to the presence of maximum degeneracy in bosonic systems. Here, we propose that a single linear dispersive sixfold excitation can be found in the electride Li12 Mg3 Si4 and its derivatives. The sixfold excitation is formed by the floating bands of elementary band representation A@12a originating from the excess electrons centered at the vacancies (i.e., the 12a Wyckoff sites). There exists a unique topological bulk-surface-edge correspondence for the spinless sixfold excitation, resulting in trivial surface “Fermi arcs” but topological hinge arcs. All gapped kz slices belong to a two-dimensional higher-order topological insulating phase, which is protected by a combined symmetry T S4z and characterized by a quantized fractional corner charge Qcorner = 3||e||/4. Consequently, the hinge arcs are obtained in the hinge spectra of the S4z-symmetric rod structure. The state with a single sixfold excitation, stabilized by both nonsymmorphic crystalline symmetries and time-reversal symmetry, is located at the phase boundary and can be driven into various topologically distinct phases by explicit breaking of symmetries, making these electrides promising platforms for the systematic studies of different topological phases.
Contribution - Problem Statement Contribution - Solution Statement Contribution - Result Statement
{ "overarching_problem_domain": "Condensed matter physics", "challenges/difficulties": "Lack of full rotational symmetry in solids leading to new excitations beyond Dirac and Weyl fermions", "research_question/goal": "Investigate sixfold excitations in electrides" } { "overarching_solution_domain": "Electrides and topological phases", "solution_approach": "Propose that a single linear dispersive sixfold excitation can be found in the electride Li12Mg3Si4 and its derivatives", "novelty_of_the_solution": "Unique topological bulk-surface-edge correspondence for the spinless sixfold excitation" } { "findings/results": "The sixfold excitation is formed by floating bands of elementary band representation A@12a. All gapped kz slices belong to two-dimensional higher-order topological insulating phase, characterized by a quantized fractional corner charge Qcorner = 3||e||/4. Hinge arcs are obtained in the hinge spectra of the S4z-symmetric rod structure.", "potential_impact_of_the_results": "Electrides are promising platforms for systematic studies of different topological phases." }
Contribution - Topic: ’Electrides’, ’Electrides in Condensed Matter Physics’, ’Higher-Order Topological Insulators’, ’Nonsymmorphic Symmetries’, ’Sixfold Excitation in Solids’, ’Sixfold Excitations’, ’Symmetry Breaking in Topological Materials’, ’Topological Bulk-Surface-Edge Correspondence’, ’Topological Phase Transitions’, ’Topological Phases in Condensed Matter Physics’, ’Topological Properties’
Title: The Tin Pest Problem as a Test of Density Functionals Using High-Throughput Calculations
Abstract: At ambient pressure tin transforms from its ground-state semi-metal $\alpha$-Sn (diamond structure) phase to the compact metallic $\beta$-Sn phase at 13 $\bullet$ C (286K). There may be a further transition to the simple hexagonal $\gamma$-Sn above 450K. These relatively low transition temperatures are due to the small energy differences between the structures, $\approx$ 20 meV/atom between $\alpha$-and $\beta$-Sn. This makes tin an exceptionally sensitive test of the accuracy of density functionals and computational methods. Here we use the high-throughput Automatic-FLOW (AFLOW) method to study the energetics of tin in multiple structures using a variety of density functionals. We look at the successes and deficiencies of each functional. As no functional is completely satisfactory, we look Hubbard U corrections and show that the Coulomb interaction can be chosen to predict the correct phase transition temperature. We also discuss the necessity of testing high-throughput calculations for convergence for systems with small energy differences.
Contribution - Problem Statement Contribution - Solution Statement Contribution - Result Statement
{ "overarching_problem_domain": "Density functionals and computational methods for phase transitions in materials.", "challenges/difficulties": "Small energy differences between phases of tin make it a sensitive test for the accuracy of density functionals.", "research_question/goal": "To study the energetics of tin in multiple structures using a variety of density functionals and assess their accuracy." } { "overarching_solution_domain": "High-throughput computational methods and density functional theory.", "solution_approach": "Using the high-throughput Automatic-FLOW (AFLOW) method to study tins energetics with various density functionals.", "novelty_of_the_solution": "Application of Hubbard U corrections to improve predictions of phase transition temperatures." } { "findings/results": "No functional is all satisfactory, but Hubbard U corrections can be chosen to predict the correct phase transition temperature.", "potential_impact_of_the_results": "Improved accuracy in predicting phase transitions in materials with small energy differences." }
Contribution - Topic: ’Convergence Testing in Computational Simulations’, ’Density Functional Theory (DFT) Accuracy’, ’High-Throughput Computational Methods’, ’Hubbard U Corrections’, ’Tin Phase Transitions’
Table 6: Examples of extracted problem/solution/result/topic contributions from scientific paper abstracts.
C.4 Distribution of Extracted Topics
This section shows the distribution of various topics extracted from the papers based on frequency. This gives us an idea of what kind of topics were extracted.
(a) Top-50 topics by frequency in decreasing order
(b) Sampled topics (every 200)
Figure 5: Distribution of topics extracted from SciPile: (a) Top-50 topics, (b) Every 200 topics. Refer §3.2 for more information.
Appendix D Compiling a Seed Hierarchy
As we discuss in §4.3, we make a few adjustments to the seed hierarchy that we obtain from Wikipedia. Specifically:
  1. 1.
    We added “Theoretical Computer Science” and “Information Theory” as separate branches under “Formal Sciences” due to their unique characteristics;
  2. 2.
    We moved “Astronomy” under “Physical Science”;
  3. 3.
    “Astronomy”, “Geology” and “Oceanography” are listed under “Earth Science” but since these topics are not specific to early, we move them up one layer so that they’re directly under “Physical Science”; The Wikipedia article groups Geology, Oceanography, and Meteorology under ;
  4. 4.
    We added “History” as a branch under “Social Sciences”;
  5. 5.
    We included “Cell Biology” and “Genetics” under “Biological Sciences” as they were relevant and their inclusion would only help in better hierarchy creation.
These modifications aim to refine the hierarchy, ensuring it accurately reflects the distinct areas within each scientific domain. The resulting hierarchy is included in Fig.6.
1{
2 "Science":{
3 "Formal Sciences":{
4 "Logic":{},
5 "Mathematics":{},
6 "Statistics":{},
7 "Computer Science":{},
8 "Information Theory":{},
9 "Systems Theory":{},
10 "Decision Theory":{}
11 },
12 "Natural Sciences":{
13 "Physical Science":{
14 "Physics":{
15 "Classical Mechanics":{},
16 "Thermodynamics and statistical mechanics":{},
17 "Electromagnetism and photonics":{},
18 "Relativity":{},
19 "Quantum Mechanics":{},
20 "Atomic and molecular physics":{},
21 "Condensed matter physics":{},
22 "Optics and acoustics":{},
23 "High energy particle physics":{},
24 "Nuclear physics":{},
25 "Cosmology":{},
26 "Interdisciplinary Physics":{}
27 },
28 "Chemistry":{
29 "Physical Chemistry":{},
30 "Organic Chemistry":{},
31 "Inorganic Chemistry":{},
32 "Analytical Chemistry":{},
33 "Biological Chemistry":{},
34 "Theoretical Chemistry":{},
35 "Interdisciplinary Chemistry":{}
36 },
37 "Earth Science":{},
38 "Astronomy":{},
39 "Geology":{},
40 "Oceanography":{},
41 "Meteorology":{}
42 },
43 "Biological Sciences":{
44 "Biochemistry":{},
45 "Cell Biology":{},
46 "Genetics":{},
47 "Ecology":{},
48 "Microbiology":{},
49 "Botany":{},
50 "Zoology":{}
51 }
52 },
53 "Social Sciences":{
54 "Anthropology":{},
55 "Economics":{},
56 "Political Science":{},
57 "Sociology":{},
58 "Psychology":{},
59 "Geography":{},
60 "History":{}
61 }
62 }
63}
Figure 6: The seed hierarchy used by our fLMSci baselines. See §D for details.
Appendix E fLMSci: LLM-based Baselines
This section includes the pipeline and prompts used for fLMSci (parallel) and fLMSci(incremental).
E.1 Pipeline for fLMSci (parallel)
This section demonstrates the pipeline used for fLMSci (par) right from extracting topics and rationales to obtaining a final taxonomy with papers. (Refer to §4.3 for more information).
Figure 7: Pipeline for of fLMSci (parallel).
E.2 Prompt for fLMSci (parallel)
This prompt guides a large language model (LLM) to expand an existing scientific taxonomy - the seed taxonomy (Refer to D) by adding a set of new topics in a logical and consistent manner. With a clear list of instructions it ensures accurate placement and also preserves the original structure. This prompt was used with Llama-3.3-70B-Instruct. (Refer to §4.3 for more information.)
Figure 8: Prompt of fLMSci (incremental) pipeline
E.3 Demonstration of actions for fLMSci (incremental)
This section demonstrates the various actions (add sibling, make parent, go down and discard) that are available for the LLM to take at various levels of taxonomy building. Refer to §4.3 for more information.
Figure 9: Actions for fLMSci (incremental)
E.4 Prompt for fLMSci (incremental)
This prompt is used to place new scientific topics into an existing seed taxonomy (Refer §D) incrementally. The model evaluates multiple possible actions based on the available action options. The prompt clearly instructs its priorities explicitly to give a hint to the model. The example usage and example output format help to get the response in a valid format. This prompt was used for Llama-3.3-70B-Instruct.
1SUBNODE_DESCRIPTIONS = {
2 "Formal Sciences": "Focuses on abstract systems and formal methodologies grounded in logic, mathematics, and symbolic reasoning. Provides theoretical frameworks (e.g., statistics, computer science, systems theory) used to model and solve problems across empirical disciplines and technology.",
3 "Natural Sciences": "Investigates the physical universe and living organisms through empirical observation, experimentation, and theoretical analysis. Includes physical sciences (e.g., physics, chemistry, astronomy) and biological sciences (e.g., genetics, ecology) to uncover fundamental laws and processes governing nature.",
4 "Social Sciences": "Studies human behavior, societies, and institutions using qualitative and quantitative methods. Encompasses disciplines like psychology, economics, and political science to analyze cultural, economic, and social interactions within historical and geographic contexts."
5}
Figure 10: Descriptive statement used for contextualizing layer-1 items in the seed hierarchy, used in fLMSci (incremental). See §4.3 for broader context.
1You are building a scientific topics based hierarchy.
2
3Path traced until now: {current_path}
4Subnode options available at this level:
5subnodes = [{subnodes}]
6New topic: "{new_topic}"
7
8Evaluate all possible actions listed below equally before choosing the most appropriate one.
9Choose the action that best preserves a logical hierarchy, semantic clarity, and appropriate abstraction level.
10
11**Priority Guidance**:
121. FIRST consider "go_down" if ANY existing subnode could reasonably contain the new topic as a specialization
132. THEN consider "make_parent" if multiple existing subnodes could be grouped under a new category
143. ONLY use "add_sibling" if the topic is FUNDAMENTALLY distinct from all existing subnodes at this level
154. "discard" should be used for low-quality or redundant topics
16
17**Critical Rules**:
18- A node about "Applications of X" should ALWAYS go under X, not as a sibling
19- Specific methods/tools belong under their parent field (e.g., "PCR" under "Molecular Biology")
20- Avoid creating flat structures
21
22Possible actions:
231) "go_down" - Use this if the topic: {new_topic} is a *more specific* subtype of one of the "subnodes" and belongs *within* it.
242) "add_sibling" - Use this if the topic: {new_topic} is on the same level of abstraction as the existing "subnodes". It should be added *alongside* them as a direct child of ‘{current_path[-1]}‘.
253) "discard" - Use this if the topic: {new_topic} is irrelevant, redundant, or already captured under another topic.
264) "make_parent" - Use this when the new topic: {new_topic} or any one of the "subnodes" is broader or more abstract than one or more of the subnodes. In that case, make the new topic a direct child of ‘{current_path[-1]}‘ and move the relevant subnodes under it. Return them in "child_nodes": [...]‘.
27
28### Example of desired usage for each action:
291) "go_down"
30 - "node": must be the name of one of the existing "subnodes"
31 - "explanation": an optional text with reasoning
32 - "child_nodes", "parent_node": not used.
33
342) "add_sibling"
35 - "node": {new_topic}
36 - "parent_node": {current_path[-1]}
37 - "explanation": optional
38 - "child_nodes": not used.
39
403) "discard"
41 - "node": {new_topic}
42 - "explanation": optional
43 - "parent_node", "child_nodes": not used
44
454) "make_parent"
46 - "node": {new_topic} or one of the "subnodes" whichever is more appropriate
47 - "child_nodes": array of the subnodes that must be moved under the new node
48 - "explanation": optional
49 - "parent_node": not used
50
51Your output must be valid JSON only:
52{{
53 "action": "go_down"|"add_sibling"|"make_parent"|"discard",
54 "node": "string",
55 "parent_node": "string or null", // only used if action = add_sibling
56 "child_nodes": ["string", ...], // only used if action = make_parent
57 "explanation": "string (optional)"
58}}
59No extra text.
Figure 11: The detailed prompt used in the execution of our fLMSci (incremental) baseline. See §4.3 for broader context.
Appendix F Further Details on Collection of Science Papers
This section provides more context on our piles of papers in our experiments from §5.1. SciPileLarge is an extension of SciPile. For each paper in SciPile, we extract five relevant keywords using an LLM (see Fig.12) and query the Semantic Scholar API [^3] https://www.semanticscholar.org/product/api with these keywords to retrieve additional relevant papers.
We apply three filtering criteria to ensure quality: (a) Citation Count: A paper must have a minimum number of citations to be considered reliable. The minimum citation count is calculated using the formula: $(2+3\times(2025-\texttt{publish\_year})$. (b) Abstract Length: A paper must have an abstract with at least 250 tokens, as measured by the tokenizer of Llama-3.1-8B-Instruct. (c) Publication Venue: A paper must be published in a recognized journal or conference. For each keyword, we select up to five papers that meet all criteria. This approach maintains the disciplinary distribution of our seed dataset SciPilewhile expanding our corpus to $10K$ papers.
Figure 12: Prompt of Keyword Extraction for Dataset Expansion
Appendix G Hyperparameters of Scychic
Here shows the models and hyperparameters we use for the experiments mentioned in §5.3. We utilize the GPT-4o model (gpt-4o-2024-08-06) to generate all contribution extractions along with detailed rationales explaining the extraction decisions. For summarizer, we use We choose Llama-3.3-70B-InstructGrattafiori et al. (2024) for its superiority of following instructions among open-source models, and use gte-Qwen2-7B-instruct as our embedder. For clustering algorithm, we apply k-means clustering. The number of clusters for each layer is (6, 40, 276) when conducting experiments on SciPile ($2K$ papers), and (6, 40, 276, 1250) when on SciPileLarge ($10k$ papers).
Appendix H Additional Experiments of Scychic
H.1 Detailed Evaluation Results on Topic Contributions
Here we show the complete evaluation results mentioned in §5.2. Scychic, fLMSci (parallel) and fLMSci(incremental) are using Topic as contribution type.
Method Accuracy (%) LLM Cost Hierarchy Structure
Strict-Acc $\uparrow$ L1-Acc $\uparrow$
Avg. # of
Input Tokens $\downarrow$
# of
Calls $\downarrow$
Max
Depth
Avg
Depth
Avg Bran.
Factor
Max Bran.
Factor
# of
Items
Contributions type: Topic
Scychic 14.9 $\pm$ 2.7 65.7 $\pm$ 4.4 5017 322 3 3 40.9 128 11k
[…]
Top-down
14.5 $\pm$ 4.7 62.5 $\pm$ 7.4 6440 322 3 3 40.9 104 11k
[…]
Bottom-up
13.9 $\pm$ 5.3 54.4 $\pm$ 12.7 3988 322 3 3 40.9 119 11k
lightgrayblack
[…]
fLMSci (par)
4.0 $\pm$ 2.8 32.0 $\pm$ 6.3 8896 226 9 6.2 13.9 734 9.9K
[…]
fLMSci (inc)
18.0 $\pm$ 5.3 91.0 $\pm$ 4.0 4040 61K 14 7.7 3.6 704 10.4K
[…]
H.2 Comparison of Different Embedding models
For the embedder mentioned in §4.1. We evaluate three embedding models—Qwen’s gte-Qwen2-7B-instruct (Li et al., 2023), OpenAI’s text-embedding-3-large, and text-embedding-ada-002. The first two performe similarly, whereas text-embedding-ada-002 produce markedly weaker results. Given the comparable overall performance between the two leading models, we selecte gte-Qwen2-7B-instruct for our main experiments due to its strong balanced performance across both metrics, superior Sctric-Acc results, and practical advantages as an open-weight model that offers greater accessibility and cost-effectiveness for reproducible research.
Models→ text-embedding-3-large gte-Qwen2-7B-instruct text-embedding-ada-002
Metrics→ L1-Acc Sctric-Acc L1-Acc Sctric-Acc L1-Acc Sctric-Acc
PROBLEM 86.7 $\pm$ 4.6 46.7 $\pm$ 0.9 81.7 $\pm$ 2.6 51.1 $\pm$ 3.8 76.0 $\pm$ 4.4 41.7 $\pm$ 5.2
SOLUTION 80.3 $\pm$ 3.4 36.7 $\pm$ 1.7 82.3 $\pm$ 1.1 48.8 $\pm$ 6.1 63.5 $\pm$ 2.0 31.0 $\pm$ 3.2
RESULTS 84.7 $\pm$ 5.7 44.0 $\pm$ 0.8 76.4 $\pm$ 6.9 46.4 $\pm$ 5.2 74.6 $\pm$ 3.4 41.0 $\pm$ 8.7
Table 8: Performance comparison across three embedding models and contribution types. gte-Qwen2-7B-instruct demonstrates superior Sctric-Acc performance across all categories, while text-embedding-3-large excels in L1-Acc for problem and results. text-embedding-ada-002 shows consistently weaker performance across both metrics.
H.3 Experiments of Prompt Engineering
We investigate the effect of different prompts on the final quality of hierarchy. In the main text, for the summarizer mentioned in §4.1, we use the detailed version prompt which is carefully curated. For comparison, we also conduct the experiments with a much simpler prompt.
Detailed (Curated) Prompt Simple Prompt
You are a scientific research expert specializing in identifying and analyzing research problems and challenges. Your task is to analyze a collection of research papers or research clusters and provide a focused analysis of the research problems they address. The input could be either a collection of individual papers or research cluster summaries. Based on the content, you need to: 1. Identify the core research problems and challenges being addressed 2. Determine the overarching problem domain that connects these research efforts 3. Analyze the specific difficulties, gaps, or limitations that motivate this research 4. Understand the research questions or goals that drive these investigations 5. Generate an appropriate cluster name that captures the essence of the problem space 6. Provide a comprehensive problem-focused analysis Here is the content to analyze: Remember to: Focus specifically on problems, challenges, and research gaps rather than solutions Be specific about the technical difficulties and limitations being addressed Identify both theoretical and practical challenges Consider interdisciplinary connections in problem formulation Maintain scientific accuracy and use precise terminology Generate only one JSON format output that must follow the structure exactly Please format your response as a JSON object with the following structure: { "Cluster Name": "A clear and specific title focusing on the problem domain (No less than 5 words)", "Problem": { "overarching problem domain": "The broad scientific domain where these problems exist", "challenges/difficulties": "Specific technical, theoretical, or practical challenges that these papers address", "research question/goal": "The fundamental research questions or objectives that motivate this work" } } You are a scientific research expert specializing in identifying and analyzing research problems and challenges. Analyze the input %s and output one JSON object: { "Cluster Name": "A clear and specific title (No less than 5 words)", "Problem": { "overarching problem domain": "", "challenges/difficulties": "", "research question/goal": "" } } Instructions Extract key themes and concepts. Identify the common thread that links the items. Craft a clear, specific title ($\geq$ 5 words) for Cluster Name. Return only the JSON—nothing else.
Table 9: Comparison of Detailed (Curated) and Simple Prompts
The results show that across all contributions, the curated prompt offers significantly better quality hierarchies.
Prompt type ↓ Embedder→ text-embedding-3-large gte-Qwen2-7B-instruct
Metrics→ L1-Acc Sctric-Acc L1-Acc Sctric-Acc
Simplified problem 75.0 $\pm$ 4.6 33.7 $\pm$ 3.7 61.0 $\pm$ 0.8 24.7 $\pm$ 1.7
Detailed 86.7 $\pm$ 4.6 46.7 $\pm$ 0.9 81.7 $\pm$ 2.6 51.1 $\pm$ 3.8
Simplified solution 65.3 $\pm$ 3.4 32.7 $\pm$ 2.6 59.0 $\pm$ 2.8 21.7 $\pm$ 2.9
Detailed 80.3 $\pm$ 3.4 36.7 $\pm$ 1.7 82.3 $\pm$ 1.1 48.8 $\pm$ 6.1
Simplified results 77.7 $\pm$ 4.1 38.0 $\pm$ 4.6 66.7 $\pm$ 3.3 27.7 $\pm$ 2.5
Detailed 84.7 $\pm$ 5.7 44.0 $\pm$ 0.8 76.4 $\pm$ 6.9 46.4 $\pm$ 5.2
Table 10: Performance comparison between simplified and detailed prompts across different embedding models and contribution types. Detailed prompts consistently outperform simplified prompts across all scenarios, with improvements ranging from 7.0 to 23.3 % for L1-Acc and 3.0 to 26.4 % for Sctric-Acc. The gte-Qwen2-7B-instruct model shows the largest performance gains, with L1-Acc improvements of 20.7, 23.3, and 9.7 % for problem, solution, and results respectively.
Appendix I Demonstration of Hierarchy
Below is a snippet of our final hierarchy result.
Figure 13: Above is a small example of a final hierarchy generated by Scychic on the SciPileLarge  dataset. The original hierarchy has 4 levels, use papers’ problem contribution. Due to space constraints, this snippet shows only two levels of clusters above the individual papers.

### v4

@@
-Scientific Literature Hierarchy Construction and Organization
+Science Hierarchography: Hierarchical Organization of Science Literature
 Abstract
@@
 Results show that our method improves interpretability and offers an alternative pathway for exploring scientific literature beyond traditional search methods.
-111 Code, data and demo are available: https://github.com/JHU-CLSP/science-hierarchography
 Science Hierarchography:
 Hierarchical Organization of Science Literature
-Muhan Gao, Jash Shah, Weiqi Wang, Daniel Khashabi Department of Computer Science Johns Hopkins University Baltimore, MD 21218, USA {mgao38,jshah48,wwang194,danielk}@jhu.edu
+Muhan Gao1, Jash Shah2, Weiqi Wang2, Kuan-Hao Huang1, Daniel Khashabi2 1Texas A&M University, 2Johns Hopkins University
 1 Introduction
@@
 2 Related Work
-Gierarchy induction: The field of taxonomy induction has progressed from early pattern-based techniques to modern LLM-augmented methods.
+Hierarchy induction: The field of taxonomy induction has progressed from early pattern-based techniques to modern LLM-augmented methods.
 Seminal work by Hearst (1992) introduced the use of hand-crafted hyponym patterns for extracting is-a relationships.
@@
 We focus on scaling taxonomy induction for the domain of scholarly literature—a setting that presents greater challenges than typical setups (e.g., entity hierarchy) due to the complexity, size, and evolving nature of scientific content.
-The most comparable effort is by Oarga et al.
-(2024), though our broader objectives require fundamentally different algorithmic strategies and operate without access to ground truth labels.
+The closest works are Oarga et al.
+(2024), which build domain-specific hierarchies (e.g., Chemical Elements) using iterative LLM refinement, and Zhu et al.
+(2025), which organizes survey-based collections of fewer than 100 papers.
+Our work differs in scope and scale: our broader objectives require fundamentally different algorithmic strategies and operate without access to ground truth labels.
 Structured representation of science: As science grows at an unprecedented rate (Teufel et al., 1999; Pertsas and Constantopoulos, 2017; Constantin et al., 2016; Fisas et al., 2016; Liakata et al., 2010), numerous frameworks have emerged to structure this information through knowledge graphs and taxonomies (Fathalla et al., 2017; Jaradeh et al., 2019; Oelen et al., 2020; Vogt et al., 2020; Soldatova and King, 2006).
@@
 In Table 1 we summarize the differences with existing hierarchical resources.
-While most prior systems are limited to one/two level(s) of depth and rely on manually assigned labels for indexing—a process often prone to bias (Hadfield, 2020)—our approach supports deeper, algorithmically generated hierarchies with semantically rich node descriptions.
+While most prior systems are limited to one/two level(s) of depth and rely on manually assigned labels for indexing—a process often prone to bias (Hadfield, 2020).
+For example, Scopus employs a fixed two-layer hierarchy (ASJC codes) assigned at the journal level, so paper-level classifications are inherited rather than content-derived.
+In contrast, our approach supports deeper, algorithmically generated hierarchies with semantically rich node descriptions.
 This enables a more flexible and interpretable representation of scientific knowledge.
+We provide a detailed comparison with existing systems in §A.
 3 Science Hierarchography: Toward Hierarchy of Scholarly Work
-We begin with a formal problem definition (§3.1), followed by content representation (§3.2) and depth considerations (§3.3).
+We begin with the problem definition (§3.1), followed by content representation (§3.2) and depth considerations (§3.3).
 3.1 Formal Problem Statement
 We define the task of Science Hierarchography as an inference problem where the input is a large set of scientific papers: $P=\{p_{1},p_{2},\ldots,p_{n}\}$.
-The goal is to infer a hierarchical structure for these papers (e.g., their problems and techniques).
-This structure represents levels of specificity and abstraction, with nodes closer to the root representing broader topics.
+The goal is to infer a hierarchical structure (i.e., a tree) for a specific contribution type (e.g., problem statement) of a collection of papers.
+The nodes of this tree are the atomic concepts representing scholarly ideas or goals.
+The edge (relations connecting two nodes) encode whether one node is a specific version of another node (i.e., “isA” relationships) which defines a hierarchical link between node pairs, indicating a child node is a subclass of its more abstract parent node (e.g., “RLHF isA RL” means “RLHF” is a type of “RL”).
+The specific papers $P$ are the nodes of this tree.
+The overall hierarchy represents levels of specificity and abstraction, with nodes closer to the root representing broader topics.
 Broader topics are at the upper levels, while more specific subtopics and individual papers are at the lower levels.
-Formally, this goal is defined as inferring a graph that encodes relationships between nodes (atomic concepts representing scholarly ideas or goals).
-This is formally defined as $T\vcentcolon=(N,E)$ where:
-  • NN is the set of all concepts (nodes) present in the hierarchy: $N=\{n_{1},n_{2},\ldots,n_{N}\}$.
-  • $E$ is the set of all edges connecting node pairs: $E=\{r_{1},r_{2},\ldots,e_{M}\}$ where each edge is a node pair $r_{i}=(n_{p},n_{c})$ connects a parent node $n_{p}$ to its child node $n_{c}$.
-  • For each edge $e\in E$, there is a type attribute: $\text{type}(e)\in\{\text{``isA''},\text{``instanceOf''}\}$ associated with the edge.
-  • Each node $n\in N$ has a property $n.prop$ which defines the content of the node (e.g., a sentence description or a few keywords).
-The edge (relations connecting two nodes) have types that is either “isA” relationships or “instanceOf.” The "isA" relationship defines a hierarchical link between node pairs, indicating a child node is a subclass of its more abstract parent node (e.g., "RLHF isA RL" means "RLHF" is a type of "RL").
-The "instanceOf" relationship indicates that a specific instance belongs to a concept category (e.g., "Ouyang et al.
-(2022) instanceOf RLHF" means "Ouyang et al.
-(2022)" is a specific instance of the broader “RLHF” cluster).
+Why a tree structure?
+A tree offers a clear, interpretable way to capture hierarchical relations among scientific ideas, showing how concepts specialize or generalize without cycles or ambiguity.
+Each node inherits meaning from its ancestors, tracing the progression from broad themes to concrete contributions.
+We build one tree per contribution type (e.g., problem, method; §3.2), forming a forest of hierarchies that naturally accommodates the multi-faceted and interdisciplinary nature of scientific research.
 3.2 Decomposing Papers to Contributions
@@
 This also mitigates the issue of input length: papers typically range from 4 to 10 pages (5K to 10K tokens), making full-document processing across large corpora infeasible and costly for LLMs.
-We use an LLM to preprocess each paper (title and abstract) and break them down into a pre-defined set of contributions, akin to prior work Hope et al.
+We use an LLM (gpt-4o-2024-08-06) to preprocess each paper (title and abstract), use the paper’s title and abstract as the input, and request the LLM to break them down into a pre-defined set of contributions, akin to prior work Hope et al.
 (2017); Chan et al.
-(2018) that mine “problem schema” from existing documents.
+(2018) that mines “problem schema” from existing documents.
 We consider the following contribution types: (1) problem statement (the problem addressed), (2) solution (the technical approach used), (3) result (the key finding), and (4) topic (the overarching themes).
@@
 We note that each contribution may include additional dimensions (sub-contributions).
-For instance, a “result” encompasses both the “outcome” and its “potential impact.”
+For instance, a “result” encompasses both the “outcome” and its “potential impact.” In total, this yields $C=11$ sub-contributions per paper.
+The LLM performs consistently during extraction: when we deliberately remove information from the input (primarily from the abstract), it correctly leaves the corresponding sub-contributions blank rather than hallucinating content, which demonstrates its reliability.
 3.3 Choosing Hierarchy Depth
@@
 In practice, we use $L=3$ for a 2K-paper corpus and $L=4$ for 10K papers, consistent with this logarithmic scaling.
-Extrapolating further, corpora of $10^{7}$ papers would likely require depths of $L=6$ or $7$.
+Extrapolating further, corpora of $10^{7}$ papers would likely require depths of $L=6$ or 77.
 4 Tackling Science Hierarchography
@@
 While all approaches leverage LLMs to some extent, they differ significantly in their reliance on them: some require many calls (linear or quadratic in the number of papers), while others are more efficient (e.g., logarithmic).
-Since our goal is to scale to millions of papers, minimizing LLM usage is critical.
-Our objective is to identify the method that yields the highest-quality hierarchy with the lowest LLM overhead, balancing quality, latency, and cost.
-4.1 Scychic: Alternating Between Clustering and Summarization
-This approach is based on the following design choices: (1) access to embedder, a neural model that converts a description into a $d$-dimensional vector, (ideally) capturing its semantic meaning; (2) a clustering algorithm clusterer that, given the hyperparameter $k$, generates $k$ clusters; (3) a contribution type (e.g., problem definition) and its dimensions $C$ extracted per paper as detailed in §3.2 which determines the focus of the node descriptions; (4) summarizer, an LLM that generates a summary description which (ideally) provides a more abstract description of a collection of node descriptions; and (5) the total number of hierarchy layers $L$ and target number of clusters in each layer $(k_{1},k_{2},\ldots,k_{L})$.
-Initialization: The approach begins by embedding each papers.
-For each paper $p_{i}$, we embed each component in $C$: $\texttt{{\color[rgb]{0.09765625,0.0078125,0.6796875}embedder}}(c_{j}^{i})\in\mathbb{R}^{d}$, where $j\in C$.
-This process results in $|C|$ embeddings per paper.
-We concatenate these embeddings, yielding $\mathbb{R}^{d.|C|}$ embeddings per paper.
-With this, we present the main algorithm which involves two phases:
-Top-down phase: First, we use a top-down strategy, iterating through the top half of the layers.
-For each layer $l\in(1\ldots\left\lfloor L/2\right\rfloor)$ (starting from the root and moving downward), we cluster the documents using clusterer.
-For the first layer ($L=1$), we partition the entire paper collection into $k_{1}$ top-level clusters using their embeddings.
-For the second level, we process each top-level cluster independently, i.e., within each top-level cluster, we take only the papers belonging to that cluster and apply clusterer to their embeddings to divide them into subgroups.
-This process continues recursively: for subsequent layers, clustering operates on the data points within each cluster.
-We run as many clustering algorithms as there are clusters in the previous layer.
-The allocation of the total cluster count across different top-level clusters is determined proportionally based on the number of papers in each top-level cluster.
-If a top-level cluster contains a larger proportion of papers, it will be allocated more subclusters from the total cluster count.
-This proportional allocation ensures that areas with higher paper density receive finer-grained clustering.
-This recursively subdivides each cluster at level $l$ into multiple clusters at level $l+1$.
-This top-down phase continues until we reach level $\left\lfloor L/2\right\rfloor$.
-Alongside clustering, we also use the summarizer to generate abstracted summaries for each of the clusters.
-The generated cluster description follows the same structure or style as the input descriptions.
-For example, if the inputs are statements about problem categories, the output from summarizer is also in the same style, but more abstract.
-Bottom-up phase: In the second phase, we switch to a bottom-up strategy to construct the remaining levels ($\left\lfloor L/2\right\rfloor+1$ through $L$).
-To form clusters for bottom-level (layer $L$), we apply clusterer to the paper embeddings within each sub-cluster within level-$\left\lfloor L/2\right\rfloor$ (the lowest level clustering obtained from top-down approach).
-We then use the summarizer to create an abstracted description for each cluster.
-We repeat this process for all layers from $L$ to $\left\lfloor L/2\right\rfloor+1$.
-To build layer $l$, we start by embedding the generated cluster summaries from the level below l1l-1 using embedder, similar to how we embedded the papers.
-We then run the clustering clusterer on these new embeddings and generate abstracted summaries for the clusters to group these summaries into higher-level clusters.
-This bottom-up aggregation continues until we connect with the previously constructed level $\left\lfloor L/2\right\rfloor$ clusters.
+Since our goal is to scale to a large number of papers, minimizing LLM usage is critical.
+Our objective is to identify a method that yields the highest-quality hierarchy with the lowest LLM overhead, balancing quality, latency, and cost.
 Algorithm 1 Scychic algorithm
 1:Set of papers $P=\{p_{1},p_{2},\ldots,p_{n}\}$, embedder, clusterer, summarizer, num of layers $L$, target cluster sizes $(k_{1},k_{2},\ldots,k_{L})$
-2:Initialization: For each paper $p_{i}\in P$, embed their selected components to form $\mathbb{R}^{d\times|C|}$.
+2:Initialization: For each paper $p_{i}\in P$, using embedder embed their selected components to form $\mathbb{R}^{d\times|C^{\prime}|}$.
 3:for layer $l=1$ to $\lfloor L/2\rfloor$ do $\triangleright$ Top-down phase
@@
 9:  Use summarizer to generate summaries for each cluster
-10:for each cluster $\kappa$ at level $\lfloor L/2\rfloor$ do $\triangleright$ Bottom-up phase
+10:for each cluster $\tau$ at level $\lfloor L/2\rfloor$ do $\triangleright$ Bottom-up phase
 11:  for layer $l=L$ to $\lfloor L/2\rfloor+1$ do
 12:   if $l=L$ then
-13:     Collect the embeddings of papers within $\kappa$.
+13:     Collect the embeddings of papers within $\tau$.
 14:   else
@@
 18:return Hierarchical structure
+4.1 Scychic: Alternating Between Clustering and Summarization
+Overview: Our method builds each contribution-type hierarchy through two complementary stages: a top-down phase that clusters paper embeddings into progressively finer subgroups and summarizes each cluster, followed by a bottom-up phase that embeds and reclusters the generated summaries to form higher-level abstractions.
+The combined process yields coherent, interpretable hierarchies that capture both fine-grained and global structure.
+Ingredients: This approach is based on the following design choices: (1) access to embedder, a neural model that converts a description into a $d$-dimensional vector, (ideally) capturing its semantic meaning; (2) a clustering algorithm clusterer that, given the hyperparameter $k$, generates $k$ clusters; (3) a contribution type (e.g., problem definition) and its dimensions $C^{\prime}$ extracted per paper as detailed in §3.2 which determines the focus of the node descriptions; (4) summarizer, an LLM that generates a summary description which (ideally) provides a more abstract description of a collection of node descriptions; and (5) the total number of hierarchy layers $L$ and target number of clusters in each layer $(k_{1},k_{2},\ldots,k_{L})$.
+Specifically, for embedder we use gte-Qwen2-7B-instruct, for our summarizer we use Llama-3.3-70B-Instruct Grattafiori et al.
+(2024), and for clusterer, we apply k-means clustering.
+(further details in §G.)
+Initialization: The approach begins by embedding each paper.
+For each paper $p_{i}$, we embed each component in $C^{\prime}$: $\texttt{{\color[rgb]{0.09765625,0.0078125,0.6796875}embedder}}(c_{j}^{i})\in\mathbb{R}^{d}$, where $j\in C^{\prime}$.
+This process results in $|C^{\prime}|$ embeddings per paper.
+We concatenate these embeddings, yielding $\mathbb{R}^{d.|C^{\prime}|}$ embeddings per paper.
+We now present the main algorithm consisting of two phases:
+Phase 1: Top-down: We begin with a top-down strategy that recursively partitions the paper set through the upper half of the hierarchy ($l\in[1,\left\lfloor L/2\right\rfloor]$).
+At the first level, all papers are clustered into $k_{1}$ groups using their embeddings.
+Each cluster is then processed independently—papers within a cluster are reclustered using clusterer to form finer subgroups.
+The number of subclusters assigned to each parent cluster scales linearly with its paper count, ensuring denser regions of the corpus receive finer resolution.
+This recursive subdivision continues until level $\left\lfloor L/2\right\rfloor$, producing a coarse-to-fine hierarchy.
+At each level, summarizer to generate abstracted summaries for each of the clusters based on the clustered papers’ titles and abstracts.
+The generated cluster description follows the same structure or style as the input descriptions.
+For example, if the inputs are statements about problem categories, the summaries are also in the same style, but more abstract.
+Phase 2: Bottom-up: In the second phase, we switch to a bottom-up strategy to construct the remaining levels ($\left\lfloor L/2\right\rfloor+1$ through $L$).
+To form clusters for bottom-level (layer $L$), we apply clusterer to the paper embeddings within each sub-cluster within level-$\left\lfloor L/2\right\rfloor$ (the lowest level clustering obtained from the top-down approach).
+We then use the summarizer to create an abstracted description for each cluster.
+We repeat this process for all layers from $L$ to $\left\lfloor L/2\right\rfloor+1$.
+To build layer $l$, we start by embedding the generated cluster summaries from the level below l1l-1 using embedder, similar to how we embedded the papers.
+We then run the clustering clusterer on these new embeddings and generate abstracted summaries for the clusters to group these summaries into higher-level clusters.
+This bottom-up aggregation continues until we connect with the previously constructed level $\left\lfloor L/2\right\rfloor$ clusters.
 Rationale behind the hybrid design: The hybrid approach merges the strengths of top-down and bottom-up strategies.
@@
 We refer to these baselines as fLMSci (pronounced “flimsy”) and present two variants below.
-Initializing a Seed Hierarchy: The first step involves creating a seed hierarchy, starting with the hierarchy of sciences from the Wikipedia page on branches of science.222en.wikipedia.org/wiki/Branches_of_science We made several adjustments to this hierarchy, detailed in §D, where the resulting seed hierarchy is also included.
-fLMSci (parallel): Adding Many Items in Parallel with Few Prompts: This approach involves adding items to the seed hierarchy in parallel.
-First, gather all unique items extracted from different papers and create batches of 100 items each (batching because all of these items would not fit within the context window of LLM).
-Then, implement a multi-threaded program where each thread adds a batch of keywords to a clone of the seed hierarchy.
-Finally, merge (with a Python program, not an LLM call) all these cloned trees into a single comprehensive structure.
-fLMSci (incremental): Expand the tree by incrementally adding items one at a time: This approach involves an iterative, layer-by-layer prompt that navigates the tree starting from the root node and performs specific actions.
-By default, the supported actions include: (a) Go down: select one of the nodes to traverse down to a lower layer in the tree; (b) Add sibling: Insert a new node at the same level; (c) Make parent: Create a new parent node; (d) Discard: Ignore the item, if no suitable location is found.
-This prompt are shown in Fig.11.
-The available actions vary depending on the tree’s current position.
-To prevent placing detailed items too high in the tree (above layer 3), we hide the actions that involve nodes (b) and (c).
-This restriction provides more control and reduces errors, as observed in pilot experiments.
-If the traversal reaches a child node, we hide the option to move further down (a).
-Another issue identified was frequent mistakes in earlier layers, likely due to broad labels where multiple categories could fit.
-To address this, we provide more context for the first layer by replacing label strings with descriptive definitions, as shown in Fig.10.
+For both methods, we use gpt-4o-2024-08-06 to extract the contributions (§3.2), and Llama-3.3-70B-Instruct to place them into the hierarchy.
+Initializing a Seed Hierarchy: The first step involves creating a seed hierarchy, starting with the hierarchy of sciences from the Wikipedia page on branches of science111en.wikipedia.org/wiki/Branches_of_science and refined through several adjustments detailed in §D.
+fLMSci (parallel): parallel addition of contributions: This approach expands the seed hierarchy in parallel using a small number of LLM calls.
+All unique contributions extracted from papers are first collected and divided into batches of 100 (to fit within the LLM’s context window).
+A multi-threaded program then assigns each batch to a separate thread, where the LLM adds those contributions to a cloned copy of the seed hierarchy.
+Finally, the cloned hierarchies are merged (via a Python script rather than additional LLM calls) into a single unified structure.
+fLMSci (incremental): Incremental tree expansion: This method builds the hierarchy iteratively by adding one contribution at a time through layer-by-layer prompting.
+Starting from the root, the model navigates the tree and performs one of four actions: (a) Go down: move to a lower-level node; (b) Add sibling: insert a new node at the same level; (c) Make parent: create a new parent node; or (d) Discard: ignore the contribution if no suitable location exists (Fig.
+11).
+Available actions depend on the current position in the tree.
+To avoid placing detailed contributions too high in the tree, we disable node-creation actions (b, c) above layer 3.
+When reaching a leaf node, the Go down action (a) is also unavailable.
+Pilot studies revealed frequent early-layer errors due to broad category labels; to mitigate this, we replaced top-level labels with descriptive definitions (Fig.
+10), improving contextual understanding and placement accuracy.
 4.4 Computational Complexity of Approaches
@@
 While our evaluation focuses on retrieval, the hierarchy’s utility extends beyond that.
-Two key design choices guide our evaluation: (a) selecting appropriate queries, and (b) choosing a reliable LLM.
-For (a), we sample paper titles and abstracts as queries.
+Our evaluation design involves two choices: (a) queries and (b) an evaluation model.
+For (a), we sample paper titles and abstracts.
 Although we considered generating language questions from papers, pilot studies showed both approaches yield similar results, so we use the simpler method.
-For (b), we use Qwen2.5-32b-instruct, which demonstrated performance closest to GPT-4 among open models in our pilot evaluations (§B).
-Starting at the top of the hierarchy, we provide the evaluator (LLM) with the query and the list of cluster descriptions (prompt in Fig.
-2).
-The model must select the cluster it deems most relevant.
-We move to the next level only if the chosen cluster contains the target paper.
-At each subsequent level, the same query is used, but only subclusters of the previously selected (and correct) cluster are shown.
-This continues until the model reaches the level containing the target paper.
-We report two metrics: Strict Accuracy (Strict-Acc), which captures how often the model selects the exact correct node, and Layer-1 Accuracy (L1-Acc), which measures how often it selects the correct top-level subtree.
+For (b), we use Qwen2.5-32b-instruct, which performed closest to GPT-4 among open models (§B.1).
+The process starts at the root: given a query and cluster descriptions (Fig.
+2), the LLM selects the most relevant cluster.
+If it contains the target paper, traversal continues recursively through subclusters until the correct paper-level node is reached.
+We report two metrics: Strict-Acc, the fraction of cases where the model finds the target node, and L1-Acc, which measures how often it correctly identifies the top-level subtree containing the target.
+Validation: We also validate the reliability of our LLM-based evaluation through both (a) human assessment and (b) evaluation based on existing human-annotated hierarchy from ORKGB.2).
 5.3 Experiment Design
-We conduct few sets of experiments to evaluate our method (Scychic, §4.1) against the baselines (§4.2, 4.3) following the proposed evaluation protocol.
-Hyperparameter details for Scychic are provided in §G.
-The experiments are structured as follows: (1) We begin by comparing all methods on “topic” contributions, the simplest contribution type (Table 5.3).
+We conduct a series of experiments to evaluate our method (Scychic, §4.1) against the baselines (§4.2, 4.3) using the proposed evaluation protocol.
+Hyperparameter settings for Scychic are detailed in §G.
+The experiments are organized as follows: (1) We first compare all methods on the simplest contribution type (“topic”) in Table 5.3.
 Due to the high computational cost, LLM-based baselines are evaluated only in this setting.
 (2) We then evaluate performance on more complex contributions (problem, solution, and results) using both SciPile and SciPileLarge to test scalability (Table 4).
-The results tables also report LLM Cost (average number of input tokens and LLM calls) and Hierarchy Structure (depth and branching factor per node).
+Each results table also reports LLM Cost (average input tokens and number of calls) and Hierarchy Structure (average depth and branching factor).
 Method Strict-Acc (%) $\uparrow$ L1-Acc (%) $\uparrow$ # of Calls $\downarrow$
@@
 Notably, these gains are achieved with a similar number of tokens and LLM calls, underscoring Scychic’s compute efficiency.
-LLM-based baselines can be far more expensive than Scychic.
+LLM-based approaches can be expensive.
 While fLMSci slightly outperforms Scychic in accuracy, it does so at the cost of a massive increase in LLM calls—making it impractical for large-scale use.
@@
 Scychic scales to larger paper corpus.
-For our scalability experiments, we evaluated Scychic on our larger $10K$ paper dataset SciPileLarge, using the problem statement contribution type.
-Due to the significant increase ($\times 5$) in corpus size, we implemented a 4-layer hierarchy instead of the 3-layer structure used previously.
+For our scalability experiments, we evaluate Scychic on our larger $10K$ paper dataset SciPileLarge, using the problem statement contribution type.
+Due to the significant increase ($\times 5$) in corpus size, we extend the hierarchy to four layers (versus three previously).
 Notably, Scychic achieved even higher L1-Acc (86.5%) on SciPileLarge compared to our smaller dataset SciPile.
@@
 The experimental results are in §H.2.
+Quality diagnostics confirm the reliability of the hierarchies.
+We further analyze cluster coherence by examining citation patterns within and across clusters.
+Out of 3,056 total citations, 2,587 (84.7%) occur between papers in the same cluster, while the remaining 469 (15.3%) are inter-cluster citations.
+The visualization and examples of inter-cluster citations can be found in §I.
 5.6 Sample Visualization of the Hierarchy
 The reader might be curious to see the resulting hierarchies.
-In §I we show a slice of the final hierarchy generated by Scychic on the SciPileLarge  dataset.
+In §J we show a slice of the final hierarchy generated by Scychic on the SciPileLarge  dataset.
 The original hierarchy has 4 levels, use papers’ problem contribution.
@@
 Ultimately, this balanced approach can foster innovation and drive progress in areas that might otherwise be overlooked, leading to a more inclusive and socially beneficial advancement of science.
-Acknowledgments
-This work is supported by ONR grant (N0001424-1-2089) and Defense Advance Research Projects Agency (DARPA) under Contract No.
-HR001125C0304.
-Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of DARPA.
-We sincerely thank Jiefu Ou and the broader JHU CLSP community for discussions and inspiration.
 Appendix A Additional Related Work
@@
 Similarly, modern tabular reasoning leverages structured representations to facilitate systematic inference and knowledge retrieval, demonstrating that such structure remains crucial (Wang et al., 2024).
+Comparison with existing hierarchical systems:
+Web of Science maintains a flat (one-level) collection of  250 research fields, which is useful for categorization.
+Given its flat structure, it is not a hierarchical structure.
+There are no parent-child relationships or summaries connecting broader and narrower concepts.
+These are best understood as only labels, not nodes in a multi-level taxonomy.
+Scopus uses a fixed-depth (2-layer) hierarchy based on research field names (ASJC Codes).
+Importantly, these codes are assigned at the journal level rather than to individual papers.
+Papers inherit classifications from their publishing journals, meaning the hierarchy is not derived from the actual paper content.
+PubMed MeSH terms provides hierarchical labeling for PubMed publications, but it functions at the level of keywords (few tokens) rather than leveraging the full richness of natural language from science papers.
+Crucially, it is organized around a fixed set of controlled terms rather than the actual semantic content of the papers, limiting its suitability for constructing dynamic or corpus-specific hierarchies.
+Additionally, because MeSH is manually curated, it introduces indexing delays—papers are only labeled after publication—and is subject to human bias, as noted by  Hadfield (2020).
+Microsoft Academic Graph (MAG), though discontinued in 2021, offered a rich graph-based structure connecting papers and authors.
+Its hierarchical classification derived primarily from citation patterns and machine learning clustering rather than semantic paper content, which limited cluster interpretability.
 Appendix B Evaluation Framework
@@
 Figure 2: Prompt used for Evaluation
+B.1 Pilot Experiment for Evaluator Choice
 One question is, which LLM should we use for evaluation?
-We chose Qwen2.5-32b-instruct for its strong instruction-following capabilities.
+As discussed in §5.2, we chose Qwen2.5-32b-instruct for its strong instruction-following capabilities.
 In pilot experiments, Qwen showed a high consistency against GPT4 score, compared to other open-weight models.
@@
 Table 5: Agreement of different evaluator LLMs against GPT4.
+B.2 Validation of LLM-based Evaluation
+As we discuss in §5.2, to validate our evaluation framework, a Computer Science PhD student analyzes 200 error cases (50 cases per layer).
+For each case, the annotator determines whether the error comes from the LLM evaluator or from the hierarchy itself.
+The analysis reveals three types of cases.
+First, only 9 cases (4.5%) are clear evaluator errors.
+Second, in 39 cases (18.5%), both the evaluator’s choice and the hierarchy path are reasonable, which is expected for interdisciplinary works.
+Third, in the remaining 152 cases (77%), the evaluator agrees with the human annotator.
+These results confirm the reliability of our LLM-based evaluation approach.
+To further validate our LLM-based evaluation approach, we downloaded the annotations from the Open Research Knowledge Graph (ORKG).
+On this website, papers are curated entirely by volunteers who are strongly familiar with the topics of the papers.
+We use the subset of the ORKG data focused on the Engineering domain.
+This led to a collection of 4.4K papers that are organized in a 2-layer hierarchy.
+Treating this data as a high-quality hierarchy, the question is whether our evaluation would assign it a high score.
+We ran our evaluation experiment with Qwen2.5-32B-Instruct as the LLM-as-a-judge.
+Similar to our setup from the paper, we use paper title/abstract as queries, and require the evaluator to traverse the hierarchy by incrementally making the most appropriate choice between all possible cluster candidates.
+Our results show that the evaluator model has an accuracy of 83% (i.e., in 83% of the runs it identified the correct paper).
+This indicates that our evaluation metric is able to assign a high score to a good hierarchy.
 Appendix C Extracting Paper Contributions
-Below are prompts and examples for extracting different contributions (problem, solution, result and topics) from papers’ titles and abstracts.
+As we discuss in §3.2, below are prompts and examples for extracting different contributions (problem, solution, result and topics) from papers’ titles and abstracts.
 we utilize the GPT-4o model (gpt-4o-2024-08-06) to generate all contribution extractions along with detailed rationales explaining the extraction decisions.
@@
 Appendix E fLMSci: LLM-based Baselines
-This section includes the pipeline and prompts used for fLMSci (parallel) and fLMSci(incremental).
+This section includes the pipeline and prompts used for fLMSci (parallel) and fLMSci(incremental) from §4.3.
 E.1 Pipeline for fLMSci (parallel)
@@
 We utilize the GPT-4o model (gpt-4o-2024-08-06) to generate all contribution extractions along with detailed rationales explaining the extraction decisions.
-For summarizer, we use We choose Llama-3.3-70B-InstructGrattafiori et al.
+For summarizer, we use Llama-3.3-70B-InstructGrattafiori et al.
 (2024) for its superiority of following instructions among open-source models, and use gte-Qwen2-7B-instruct as our embedder.
@@
 The gte-Qwen2-7B-instruct model shows the largest performance gains, with L1-Acc improvements of 20.7, 23.3, and 9.7 % for problem, solution, and results respectively.
-Appendix I Demonstration of Hierarchy
-Below is a snippet of our final hierarchy result.
+Appendix I Visualization and Examples of Inter-Cluster Citations
+Below is the figure of comparing inter- and intra-cluster citation counts mentioned in §$5.$
+Figure 13: Comparison of inter(green)- and intra(blue)-cluster citation counts.
+84.7% citations are between papers in the same top layer cluster, and the rest inter-cluster citations are mostly theory-to-application works, which proves the reliability of Scychic.
+Below are examples of citations between papers in different top-layer clusters.
+These examples show that many inter-cluster citations represent theory-to-application connections, while the last row illustrates cross-disciplinary citations between research fields.
+Importantly, all papers involved are correctly categorized—the inter-cluster citations reflect legitimate relationships rather than classification errors.
+Citing Paper Cited Paper
+Rationale: AI research grounding in foundational cognitive science theory
+[1pt/2pt]
+[…]
+Rationale: Theory-to-application for THz photonics
+[1pt/2pt]
+[…]
+Rationale: Hardware implementation citing quantum network theory
+[1pt/2pt]
+[…]
+Rationale: Manufacturing citing characterization techniques
+[1pt/2pt]
+[…]
+Rationale: Cross-disciplinary bridge between biology and quantum physics
+[1pt/2pt]
+[…]
+Table 11: Examples of cross-cluster citations. Each row shows the citing paper, the cited paper, their cluster names, and the citation rationale.
+Appendix J Demonstration of Hierarchy
+Below is a snippet of our final hierarchy result as mentioned in §5.6.
 Figure 14: Above is a small example of a final hierarchy generated by Scychic on the SciPileLarge  dataset.